# Texture-aware and Shape-guided Transformer for Sequential DeepFake Detection

### v2

@@ 3.2 Sequential Decoder
${\rm MCA}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{2})={\rm Concat}(h_{1},...,h_{n})W^{o},$
@@
2) Bidirectional Interactive Cross-attention. Let $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$ represent the features from the coarse-grained and fine-grained branches. Let $\mathbf{x}^{s}$ be the embedding of sequential manipulation annotations. In our task, it is intuitive to perform separate MCA operations on $\mathbf{x}^{s}$ with $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$. However, this integration fails to explore deeper connections between these three features. Inspired by the concept of multi-modal cross-attentions Yu et al. (2019); Wei et al. (2020); Zhang et al. (2023), we propose Bidirectional Interactive Cross-attention (BICA) which employs a multi-level fusion strategy to fully excavate the underlying correlations. The overview of this module is shown in Fig. 3.
Figure 3: Overview of the sequential decoder (top). Illustration of Bidirectional Interactive Cross-attention (BICA). MCA denotes Multi-head Cross-attention.
Firstly, we perform MSA on $\mathbf{x}^{s}$ to find the correlations of sequential manipulation annotations, which is then subjected to MCA with $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$ to build the cross-attention. Then these processed features are summed up and applied LayerNorm to generate multi-modal features $\mathbf{a}^{c}$ and $\mathbf{a}^{f}$. These operations can be written as
$\tilde{\mathbf{x}}^{s}$ $={\rm MSA}(\mathbf{x}^{s}),$ (9)
$\tilde{\mathbf{a}}^{c}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c}),\mathbf{a}^{c}={\rm LN}(\tilde{\mathbf{a}}^{c}+\tilde{\mathbf{x}}^{s}),$
$\tilde{\mathbf{a}}^{f}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f}),\mathbf{a}^{f}={\rm LN}(\tilde{\mathbf{a}}^{f}+\tilde{\mathbf{x}}^{s}).$
We then perform interactive cross-attention between features of $\mathbf{a}^{c}$ and $\mathbf{a}^{f}$ in two directions as
$\tilde{\mathbf{a}}^{c,f}$ $={\rm MCA}(\mathbf{a}^{c},\mathbf{a}^{f},\mathbf{a}^{f}),\mathbf{a}^{c,f}={\rm LN}(\tilde{\mathbf{a}}^{c,f}+\mathbf{a}^{c}),$ (10)
$\tilde{\mathbf{a}}^{f,c}$ $={\rm MCA}(\mathbf{a}^{f},\mathbf{a}^{c},\mathbf{a}^{c}),\mathbf{a}^{f,c}={\rm LN}(\tilde{\mathbf{a}}^{f,c}+\mathbf{a}^{f}).$
Next, we fuse the sequential features $\tilde{\mathbf{x}}^{s}$ with the enhanced spatial features $\mathbf{a}^{c,f}$ and $\mathbf{a}^{c,f}$, following the operations as
$\mathbf{A}$ $=[\mathbf{a}^{c},\mathbf{a}^{f},\mathbf{a}^{c,f},\mathbf{a}^{f,c}],$ (11)
$\tilde{\mathbf{x}}^{o}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{A},\mathbf{A}),$
${\mathbf{x}}^{o}$ $={\rm LN}(\tilde{\mathbf{x}}^{o}+\tilde{\mathbf{x}}^{s}).$
Through this fusion strategy, the sequential and spatial features can be fully interacted with and expose the underlying correlations.
Shape-guided Gaussian Mapping In comparison to general image-sequence tasks, sequential manipulation annotations in our case are relatively short and may not provide sufficient guidance. To address this limitation, we propose a Shape-guided Gaussian Mapping strategy to leverage the strong spatial position priors associated with each manipulation component in the faces. Specifically, we assume that the position of each facial manipulation follows a Gaussian distribution. By learning these distributions, we can determine the probability of each position on a face image being manipulated. To simulate these Gaussian distributions, we construct a Variational Auto-encoder (VAE) Kingma and Welling (2013) with the image $\mathbf{I}$ as input. The VAE architecture contains an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder transforms the input feature into means $\bm{\mu}$ and standards $\bm{\sigma}$ of these Gaussian distributions, while the decoder reconstructs the input images from the sampled latent code. Through the reconstruction process, the encoder learns to capture the spatial manipulation information and formulate them as Gaussian distributions. Next, we create a grid map $M$ that is composed of the $x$- and $y$- coordinates of each position. This grid map is then used to predict the probability of each position in different facial manipulations. These operations can be formulated as
$\bm{\mu},\bm{\sigma}=\mathcal{E}(\mathbf{I}),$ (12)
$S=\mathcal{N}(M\;|\;\bm{\mu},\bm{\sigma}),$
where $S$ is the probability map for each spatial position. This map is integrated into the MCA operations ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})$ and ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f})$ in Eq. (9) respectively. We take the integration in ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})$ for example as
$Q_{i}=\tilde{\mathbf{x}}^{s}W^{Q}_{i},K_{i}=\mathbf{x}^{c}W^{K}_{i},V_{i}=\mathbf{x}^{c}W^{V}_{i},$ (13)
$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}}+\eta\cdot S)V_{i},$
${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})={\rm Concat}(h_{1},...,h_{n})W^{o}.$
By inserting the shape priors of facial manipulation, the cross-attention is greatly enhanced to capture the spatial and sequential forgery traces.
Figure 4: The impact of the sequential manipulation on facial attributes. The second manipulation impacts the first manipulation on (left) eye and (right) eyeglasses.
Inverted Order Prediction. Based on our analysis, we invert the annotation of the manipulation sequence for training. For instance, given a manipulated face with annotation of “eyes-nose-eyebrow-lips-hair”, we reorder it as “hair-lips-eyebrow-nose-eyes”, and convert it to $\mathbf{x}^{s}$ for sequential modeling. By doing so, the model first predicts the last manipulation of “hair”. By isolating the disturbance caused by “hair”, the model can then focus on predicting the next manipulation, which is “lips”. The rationale behind this strategy is that the manipulation of the nose can affect the manipulation of the eyes, making the detection of eye manipulations more challenging. On the other hand, the manipulation in hair is not influenced by other manipulations, making it more easily detectable (see Fig. 4). The advantage of this strategy is that, compared to forward sequence predictions, each step in the inverted sequence predictions is only influenced by itself and is independent of other types of manipulations.
3.3 Loss Function
In the training phase, we employ an auto-regressive mechanism Graves (2013) to model sequence relationships with the cross-entropy loss $L_{\rm ce}$. This loss measures the error of predicted manipulation sequence and ground truth. For the shape-guided module, we perform the general image reconstruction loss $L_{\rm rec}$ and KLD loss $L_{\rm kld}$ used in VAE Kingma and Welling (2013). The overall loss function is expressed as
$L=L_{\rm ce}+\lambda_{1}L_{\rm rec}+\lambda_{2}L_{\rm kld},$ (14)
@@ 3.3 Loss Function
Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-Acc
Multi-Cls (ResNet-34) 69.66 50.54 66.99 46.68 68.33 48.61
Multi-Cls (ResNet-50) 69.65 50.57 66.66 46.00 68.16 48.29
DETR (ResNet-34) 69.87 50.63 67.93 48.15 68.90 49.39
DETR (ResNet-50) 69.75 49.84 67.62 47.99 68.69 48.92
DRN (ResNet-34) 66.06 45.79 64.42 43.20 65.24 44.50
MA (ResNet-34) 71.31 52.94 67.58 47.48 66.45 50.21
Two-Stream (ResNet-34) 71.92 53.89 66.77 46.38 69.35 50.14
SeqFake-Former (ResNet-34) 70.88 52.96 67.85 48.36 69.37 50.66
SeqFake-Former (ResNet-50) 69.41 50.22 68.38 48.97 68.90 49.60
TS-Former (Ours) (ResNet-34) 74.70 58.43 69.13 49.68 71.92 54.06
TS-Former (Ours) (ResNet-50) 73.32 56.08 68.86 49.67 71.09 52.88
Table 2: Ablation study on the facial-components datasets.
Setting Resnet-34 Resnet-50
Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-ACC
Baseline 69.65 (+0.00) 50.43 (+0.00) 68.84 (+0.00) 49.40 (+0.00)
Baseline+DPDA 70.24 (+0.59) 51.87 (+1.44) 69.23 (+0.39) 51.05 (+1.65)
Baseline+DPDA+BICA 71.00 (+1.35) 52.67 (+2.24) 71.55 (+2.71) 53.06 (+3.66)
Baseline+DPDA+BICA+IOP 73.32 (+3.67) 56.65 (+6.22) 72.65 (+3.81) 55.03 (+5.63)
Baseline+DPDA+BICA+IOP+SGM 74.70 (+5.05) 58.43 (+8.00) 73.32 (+4.48) 56.08 (+6.68)
@@ 4.1 Experimental Settings
Datasets. Our experiments are conducted on Sequential DeepFake Dataset (SDD) Shao et al. (2022). This dataset contains two types of sequential DeepFake manipulations, the facial components manipulation and the facial attributes manipulation. The first track comprises $35,166$ face images that are subjected to replacing operations on specific facial components. There are $28$ manipulation types in this track, with a maximum sequence length of five. The second track contains $49,920$ face images with $26$ types of attribute manipulations. The maximum length of each manipulation is also five. Its split of training, validating, and testing set is $8:1:1$.
Evaluation Metrics. In line with previous work Shao et al. (2022), we employ Fixed Accuracy (Fixed-Acc) and Adaptive Accuracy (Adaptive-Acc) for evaluation. To calculate the Fixed-Acc score, the sequence annotations shorter than five are padded with the label “no manipulation” and then compare the predictions with the padded sequence annotation. On the other hand, due to the nature of the auto-regress mechanism, the model can stop predicting upon encountering the End-of-Sequence (EOS) symbol, resulting in various lengths of the predicted sequence. The adaptive-Acc score is calculated by comparing the predictions with the sequence annotations in terms of dynamic lengths.
Implementation Details. We utilize ResNet-34 He et al. (2015) as our CNN stem in main experiment. Each branch in the spatial encoder has two Transformer blocks. Each block contains four attention heads. The sequential decoder consists of two Transformer blocks. In the training phase, the image size is set to $256\times 256$ with augmentations such as horizontal Flipping and adjusting the brightness of the image. We employ the AdamW optimizer with a weight decay of $10^{-4}$. The initial learning rates are set to $10^{-4}$ for CNN stem and $10^{-4}$ for Transformer architecture. We utilize a warm-up strategy of $20$ epochs, with a batch size of $32$. The total epochs are set to $170$, with a $10\%$ reduction in learning rate every $50$ epochs. The parameters in the loss function are set as: $\lambda_{1}=0.1,\lambda_{2}=0.01$.

### v4

@@ 3.2 Multi-source Cross-attention (MSCA)
${\rm MCA}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{2})={\rm Concat}(h_{1},...,h_{n})W^{o}$
@@
Operations of MSCA. Let $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$ represent the features from the coarse-grained and fine-grained branches. Let $\mathbf{x}^{s}$ be the embedding of inverted sequential manipulation annotations. In our task, it is intuitive to perform separate MCA operations on $\mathbf{x}^{s}$ with $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$. However, this integration fails to explore deeper connections between these three features. As such, we propose a Multi-source Cross-attention (MSCA) module to capture the correlations among multiple sources. The overview of the entire module is illustrated in Fig. 4.
Firstly, we concatenate coarse-grained features $\mathbf{x}^{c}$ and fine-grained features $\mathbf{x}^{f}$ along the channel dimension. Subsequently, we facilitate interactions between the two different levels of fabricated features throuth an MLP structure, yielding shallowly fused sequencetial features $\mathbf{x}^{m}$. Then, by performing self-attention computations on $\mathbf{x}^{s}$, we extract features that exploit correlations among annotations of sequence tampering. These operations can be formulated as follows:
$\small\mathbf{x}^{m}={\rm MLP}({\rm Concat}(\mathbf{x}^{c},\mathbf{x}^{f})),\;\tilde{\mathbf{x}}^{s}={\rm MSA}(\mathbf{x}^{s})$ (8)
Next, we implement a second-stage fusion strategy, performing multi-head cross-attention between $\tilde{\mathbf{x}}^{s}$ and $\mathbf{x}^{c}$, $\mathbf{x}^{m}$, $\mathbf{x}^{f}$ respectively, as:
$\tilde{\mathbf{a}}^{c}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c}),\mathbf{a}^{c}={\rm LN}(\tilde{\mathbf{a}}^{c}+\tilde{\mathbf{x}}^{s}),$ (9)
$\tilde{\mathbf{a}}^{m}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{m},\mathbf{x}^{m}),\mathbf{a}^{m}={\rm LN}(\tilde{\mathbf{a}}^{m}+\tilde{\mathbf{x}}^{s}),$
$\tilde{\mathbf{a}}^{f}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f}),\mathbf{a}^{f}={\rm LN}(\tilde{\mathbf{a}}^{f}+\tilde{\mathbf{x}}^{s})$
Finally, we directly sum up the three types of features and then pass them through an MLP to filter out irrelevant noise, thereby uncovering identical manipulation patterns and obtaining the final sequential deepfake features, following the operations as
$\small\mathbf{A}=\tilde{\mathbf{a}}^{c}+\tilde{\mathbf{a}}^{m}+\tilde{\mathbf{a}}^{f},\;\tilde{\mathbf{x}}^{o}={\rm MLP}(A),\;\mathbf{x}^{o}={\rm LN}(\tilde{\mathbf{x}}^{o}+\mathbf{A})$ (10)
Figure 5: Integration with MSCA.
@@
FACC AACC FACC AACC FACC AACC
Multi-Cls (ResNet-34) [28] 69.66 50.54 66.99 46.68 68.33 48.61
Multi-Cls (ResNet-50) [28] 69.65 50.57 66.66 46.00 68.16 48.29
DETR (ResNet-34) [2] 69.87 50.63 67.93 48.15 68.90 49.39
DETR (ResNet-50) [2] 69.75 49.84 67.62 47.99 68.69 48.92
DRN [29] 66.06 45.79 64.42 43.20 65.24 44.50
DRN [29] 69.38 49.76 67.66 47.31 68.52 48.54
MA [39] 71.31 52.94 67.58 47.48 69.45 50.21
MA [39] 70.12 51.82 68.91 49.19 69.52 50.51
Two-Stream [18] 71.92 53.89 66.77 46.38 69.35 50.14
Two-Stream [18] 72.00 53.97 65.20 44.39 68.60 49.18
SeqFake-Former (ResNet-34) [23] 72.13 54.80 67.99 48.32 70.06 51.56
SeqFake-Former (ResNet-34) [23] 70.97 52.65 68.40 48.93 69.69 50.79
SeqFake-Former (ResNet-50) [23] 72.65 55.30 68.86 49.63 70.76 52.47
SeqFake-Former (ResNet-50) [23] 69.12 50.03 67.93 48.65 68.53 49.34
MMNet (ResNet-50) [31] 73.93 56.83 69.27 50.44 71.60 53.64
TSOM (ResNet-34) 75.21 59.74 69.18 50.54 72.20 55.14
TSOM (ResNet-50) 75.53 59.67 69.80 51.21 72.67 55.44
Table 2: Performance of different methods on challenging scenarios.
Methods Facial-components Facial-attributes Average
FACC AACC FACC AACC FACC AACC
Two-Stream 24.00 12.74 31.55 16.92 27.78 14.83
DRN 48.11 14.11 50.35 20.62 49.23 17.37
MA 59.08 33.38 67.20 47.07 63.14 40.23
SeqFake-Former (ResNet-34) 58.70 31.43 62.88 41.88 60.79 36.66
SeqFake-Former (ResNet-50) 57.35 29.50 64.81 44.49 61.08 37.00
TSOM (ResNet-34) 61.79 36.80 67.44 47.14 64.62 41.97
TSOM (ResNet-50) 62.52 37.21 67.33 48.03 64.93 42.62
3.3 Shape-guided Gaussian Mapping
We propose a Shape-guided Gaussian Mapping (SGM) strategy to leverage the strong spatial position priors associated with each manipulation component in the faces. Since each manipulation has its specific characteristic, its manipulation shape likely follows a certain distribution. By learning these distributions, we can predict the probability of each position on a face image being manipulated. This probability can serve as shape priors to further enhance the effectiveness of MSCA.
Distribution Estimation. Since no ground truth is provided, estimating these distributions is non-trivial. Inspired by unsupervised learning strategies, we construct a Variational Auto-encoder (VAE) [11] to estimate these distributions. Specifically, we construct an encoder $\mathcal{E}$, which transforms the input image $\bm{\mathcal{I}}$ into means $\bm{\mu}$ and standards $\bm{\sigma}$ of these distributions as $\bm{\mu},\bm{\sigma}=\mathcal{E}(\bm{\mathcal{I}})$. Note that $\bm{\mu},\bm{\sigma}$ are vectors containing mean and standard of each manipulation. To ensure these distributions align with the input image, we employ a decoder $\mathcal{D}$ for reconstruction based only on the estimated distributions.
Integration with MSCA. We create a grid map $M$ that is composed of the $x$- and $y$- coordinates of each position. This grid map is then used to predict the probability of each position in different facial manipulations. This process can be formulated as $S=\mathcal{N}(M\;|\;\bm{\mu},\bm{\sigma})$, where $S$ is the probability map for each spatial position. Considering different heads in MSCA have different focuses. To properly integrate SGM with MSCA, we craft different probability maps $\{S^{\prime}_{i}\}_{i=1}^{h}$ for each head. Specifically, we use $\bm{\mu},\bm{\sigma}$ as the base of all heads and predict the head-specific offset $\Delta\bm{\mu}_{i},\Delta\bm{\sigma}_{i}$ for $i$-th head. These offsets are obtained as $\Delta\bm{\mu}_{i},\Delta\bm{\sigma}_{i}={\rm FC}(\mathcal{E}_{f}(\bm{\mathcal{I}}))$, where $\mathcal{E}_{f}(\bm{\mathcal{I}})$ denotes the last feature of encoder $\mathcal{E}$. Therefore, the probability map $S^{\prime}_{i}$ for $i$-th head can be defined as
$\small S^{\prime}_{i}=\mathcal{N}(M\;|\;\bm{\mu}+\Delta\bm{\mu_{i}},\bm{\sigma}+\Delta\bm{\sigma_{i}})$ (11)
This process is illustrated in Fig. 5. Then these maps are integrated into ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})$ and ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f})$ operations in MSCA, by calculating the attention as
$\small h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}}+\eta S^{\prime}_{i})V_{i}$ (12)
3.4 Inverted Order Prediction
In the training phase, we invert the annotation of the manipulation sequence. For instance, as shown in Fig. 4, the manipulated face is annotated as “eye-nose-eyebrow-hair-lip”, we rearrange it to ”lip-hair-eyebrow-nose-eye” as the input of the MSCA module for sequential modeling. This strategy enables the model to first predict the final manipulation (e.g., hair) and understand the consequent disturbance caused by it. Then the model concentrates on the preceding manipulation (e.g., lips). The rationale behind this strategy is that the subsequent manipulation (e.g., noise) may affect the preceding ones (e.g., eye), making it more challenging to detect them (recall Fig. 2). In contrast, preceding manipulation (e.g., hair) is not influenced by subsequent ones, making it more easily detectable. In summary, this strategy offers the advantage that, compared to forward sequence predictions, each step in the inverted sequence predictions is only influenced by itself and remains independent of other manipulations.
3.5 Objectives
In the training phase, the auto-regressive mechanism [5] is employed to model sequence relationships with the cross-entropy loss $L_{\rm ce}$. This loss measures the error of the predicted manipulation sequence and ground truth. To learn the shape priors, we perform the image reconstruction loss $L_{\rm rec}$ and KLD loss $L_{\rm kld}$ on constructed auto-encoder $(\mathcal{E},\mathcal{D})$ as in [11].
The overall loss function is expressed as
$L=L_{\rm ce}+\lambda_{1}L_{\rm rec}+\lambda_{2}L_{\rm kld}.$ (13)
Table 3: Effect of each component.
Setting Resnet-34 Resnet-50 Average
FACC AACC FACC AACC FACC AACC
BL 70.45 (+0.00) 52.18 (+0.00) 69.47 (+0.00) 50.89 (+0.00) 69.96 (+0.00) 51.34 (+0.00)
BL+DPDA 71.07 (+0.62) 52.84 (+0.66) 71.06 (+1.59) 52.80 (+1.91) 71.07 (+1.11) 52.82 (+1.48)
BL+DPDA+MSCA 71.74 (+1.29) 54.05 (+1.87) 72.04 (+2.57) 54.42 (+3.53) 71.89 (+1.93) 54.24 (+2.90)
BL+DPDA+MSCA+IOP 74.22 (+3.77) 58.19 (+6.01) 74.30 (+4.83) 58.05 (+7.16) 74.26 (+4.30) 58.12 (+6.78)
BL+DPDA+MSCA+IOP+SGM 75.21 (+4.76) 59.74 (+7.56) 75.53 (+6.06) 59.67 (+8.78) 75.37 (+5.41) 59.71 (+8.37)
Table 4: Performance of applying IOP on various methods.
Methods Facial-components Facial-attributes Average
FACC AACC FACC AACC FACC AACC
Two-Stream 72.00 53.97 65.20 44.39 68.60 49.18
Two-Sream+IOP 73.55(+1.55) 58.03(+4.06) 67.60(+2.40) 47.78(+3.39) 70.58(+1.98) 52.91(+3.73)
DRN 69.38 49.76 67.66 47.31 68.52 48.54
DRN+IOP 71.81(+2.43) 54.21(+4.45) 68.89(+1.23) 49.74(+2.43) 70.35(+1.83) 51.98(+3.44)
MA 70.12 51.82 68.91 49.19 69.52 50.51
MA+IOP 73.41(+3.29) 57.30(+5.48) 68.13(-0.78) 48.58(-0.61) 70.77(+1.25) 52.94(+2.43)
SeqFake-Former 70.97 52.65 68.40 48.93 69.69 50.79
SeqFake-Former+IOP 73.34 (+2.37) 56.93 (+4.28) 69.09(+0.69) 48.40 (-0.53) 71.22 (+1.53) 52.67(+1.88)
@@ 4.1 Experimental Settings
Datasets. To the best of our knowledge, there is only one public dataset for sequential DeepFake detection that proposed in [23]. This dataset contains two types of sequential DeepFake manipulations, the facial components manipulation and the facial attributes manipulation. The first track comprises $35,166$ face images that are subjected to replacing operations on specific facial components. There are $28$ manipulation types in this track, with a maximum sequence length of five. The second track contains $49,920$ face images with $26$ types of attribute manipulations. The maximum length of each manipulation is also five. Its split of training, validating, and testing set is $8:1:1$.
Evaluation Metrics. In line with previous work [23], we employ Fixed Accuracy (FACC) and Adaptive Accuracy (AACC) for evaluation. To calculate the FACC score, the sequence annotations shorter than five are padded with the label “no manipulation” and then compare the predictions with the padded sequence annotation. On the other hand, due to the nature of the auto-regress mechanism, the model can stop predicting upon encountering the End-of-Sequence (EOS) symbol, resulting in various lengths of the predicted sequence. The AACC score is calculated by comparing the predictions with the sequence annotations in terms of dynamic lengths.
Implementation Details. Each branch in the spatial encoder has two Transformer blocks. Each block contains four attention heads. The sequential decoder consists of two Transformer blocks. In the training phase, the image size is set to $256\times 256$ with augmentations such as horizontal Flipping and adjusting the brightness of the image. We employ the AdamW optimizer with a weight decay of $10^{-4}$. The initial learning rates are set to $10^{-4}$ for CNN stem and $10^{-4}$ for Transformer architecture. We utilize a warm-up strategy of $20$ epochs, with a batch size of $32$. The total epochs are set to $170$, with a $10\%$ reduction in learning rate every $50$ epochs. The parameters in the loss function are set as: $\lambda_{1}=0.1,\lambda_{2}=0.01$.